OpenStudy (anonymous):

Log the coordinates of the specific points in space to which your spacecraft will travel. Please remember to include the graph of your points and the lines connecting each point along with your work. Launch Area:___(1, 2)___ Point A:__________ Point B:__________ Point C:__________ You must show your work on each question below.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Idk, are you supposed to pick points?

OpenStudy (anonymous):

well i mean, is there certain points i have to pick? & theres more to the problem i just broke it down.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

"...which YOUR spacecraft will travel..." seems to imply there were some questions earlier to this question dealing with equation for the flight path.

OpenStudy (anonymous):

so could i do (2,3) (4,5) & (3,2)?

OpenStudy (ranga):

Were there other parts to this question before this one?

OpenStudy (anonymous):

Now it is your turn to plan the trajectories required to launch a spacecraft through a specific route in space. The launch area is identified on the map below. Select three points for your spacecraft to travel through and label them Point A, Point B, and Point C. Be sure that your points are not all in the same line.

OpenStudy (ranga):

Choose three points not on the same line. (1,2) is given, You have chosen: (2,3) (4,5) & (3,2). (1,2), (2,3) and (4,5) will all be on the same line.

OpenStudy (anonymous):

so (3,2) (5,5) & (6,1)

OpenStudy (ranga):

Instead of y going up and down I will just choose three points with both x and y increasing.

OpenStudy (anonymous):

im confused..

OpenStudy (ranga):

If you look at your 3 y values, it goes from2 to 5 to 1. How about (2,3), (3,5), (4, 9)? The numbers are easy to plot on a graph paper without using large numbers for x and y and using a scale factor.

OpenStudy (anonymous):

how about 1 & 2? 1.Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. 2.Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B.

OpenStudy (anonymous):

i have 8 questions. @ranga

OpenStudy (anonymous):

@phi

OpenStudy (ranga):

Launch point (1,2) A is (2,3); B is (3,5); C is (4,9) What is the equation of the line connecting (1,2) and (2,3)? Find the slope first.

OpenStudy (phi):

Have you picked your points ?

OpenStudy (anonymous):

2-3/1-2?

OpenStudy (ranga):

Yes, but use parenthesis.

OpenStudy (anonymous):

-1/-1 = 1?

OpenStudy (ranga):

Yes. Can you find the equation of the line with slope 1 and passing through any one of the above two points?

OpenStudy (anonymous):

which point do you want me to use?

OpenStudy (ranga):

first one is simpler. Either one will give you the same answer.

OpenStudy (anonymous):

y-2=1(x-1)?

OpenStudy (ranga):

Yes. Simplify it first and then put it in the standard form: Ax + By = C

OpenStudy (anonymous):

so would it be y=1x+1 ?

OpenStudy (phi):

that is in slope-intercept form... correct, but they want standard form

OpenStudy (ranga):

You can drop the 1 in front of the x. y = x + 1. But this is the slope-intercept form. They are asking for standard form. -x + y = 1 or x - y = -1

OpenStudy (anonymous):

so how would i change? i forget.

OpenStudy (anonymous):

so x-y=-1?

OpenStudy (ranga):

some text books prefer coefficient of x to be positive. But I think either one should be okay: -x + y = 1 or x - y = -1.

OpenStudy (anonymous):

#2?

OpenStudy (anonymous):

so would i use (3,5) & (4,9)?

OpenStudy (ranga):

Find the equation of the line passing through (2,3) and (3,5) just like before but this time leave the answer in slope-intercept form.

OpenStudy (anonymous):

3-5/2-3?

OpenStudy (ranga):

once again, use parenthesis.

OpenStudy (anonymous):

-2/-1 = 2?

OpenStudy (ranga):

Yes. Finish the rest.

OpenStudy (anonymous):

y-3=2(x-2) ?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

y=2x-1?

OpenStudy (ranga):

yes. That is the slope-intercept form of that line.

OpenStudy (anonymous):

ok 3& 4.

OpenStudy (anonymous):

3.Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. 4.In question 2, you selected one of two points (Point A or Point B) to be included in your point-slope equation. Write the point-slope form of that equation again, using the other point’s coordinates.

OpenStudy (ranga):

Same procedure as the previous one except here the points are (3,5) and (4,9).

OpenStudy (anonymous):

(5-9)/(3-4)?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

-4/-1 = 4?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

y-5=4(x-3)?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

y=4x-7

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

btw, thank you so much for your help!

OpenStudy (ranga):

#4 refers back to #2 where we found the equation of the line passing through (2,3 and (3,5) You found the slope m as (3-5)/(2-3) = -2/-1 = 2. Then you used the first point (2,3) to find the equation. Here they are asking you to use the second point. What is the equation of the line with slope 2 passing through (3,5)?

OpenStudy (anonymous):

y-5=2(x-3)?

OpenStudy (ranga):

yes. simplify.

OpenStudy (anonymous):

y=2x-1

OpenStudy (ranga):

Yes. It should be the same answer as you got for #2.

OpenStudy (anonymous):

#5&6?

OpenStudy (anonymous):

5.Convert the equations you arrived at in question 2 and question 4 into slope-intercept form. 6.Does the point you select matter when your write a point-slope equation? Explain your reasoning using complete sentences.

OpenStudy (ranga):

Oops. Just realized for #2 and #4 they were asking for point-slope form and NOT slope-intercept form. So for #2 leave the answer at: y - 3 = 2(x - 2). It was not necessary to go the next step. Similarly for #4, leave the answer at y - 5 = 2(x - 5).

OpenStudy (anonymous):

you mean y-5=2(x-3)?

OpenStudy (anonymous):

for # 4?

OpenStudy (ranga):

yes, that was a typo. y - 5 = 2(x-3)

OpenStudy (ranga):

Now for #5, do the next step and put it in point-slope form: From #2: y - 3 = 2(x - 2) y - 3 = 2x - 4 y = 2x - 1 From #4: y - 5 = 2(x - 3) y - 5 = 2x - 6 y = 2x - 1

OpenStudy (anonymous):

#6?

OpenStudy (ranga):

#6. The point you select does NOT matter when your write a point-slope equation. You can select either point and you will get the same answer. That is because we initially used BOTH point to calculate the slope. After that it does not matter which point is chosen for the point-slope form because the slope and either of the points will ensure that the line passes through BOTH points.

OpenStudy (anonymous):

wait im confused..

OpenStudy (ranga):

We started with two points and first found the slope. Then we used the slope and one of the points to find the equation of the line in point-slope form. They are asking if it matters which point is chosen. That is, y - y1 = m(x-x1) or should it be y - y2 = m(x - x2) The answer is it does not matter. Both will simplify to the same equation.

OpenStudy (ranga):

You can put m = (y2-y1)/(x2-x1) into the above two equations and simplify and prove mathematically they are one and the same. But I don't think they are asking for mathematical proof. Just explanation. And that is what I did 2 replies ago.

OpenStudy (anonymous):

7&8? 7.Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences. 8.Explain, using complete sentences, why it is important to understand any limitations on the domain and range.

OpenStudy (anonymous):

#6 i put "It does not matter which way you write it.. y-y1=m(x-x1) or y-y2=m(x-x2) either way it comes out the same. They will both simplify to the same equation."

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

#7?

OpenStudy (ranga):

But include that when simplified by replacing m with (y2-y1)/(x2-x1) both will yield the same result.

OpenStudy (anonymous):

what?

OpenStudy (ranga):

#6 i put "It does not matter which way you write it.. y-y1=m(x-x1) or y-y2=m(x-x2) either way it comes out the same. Since m = (y2-y1)/(x2-x1) they will both simplify to the same equation."

OpenStudy (anonymous):

#8?

OpenStudy (ranga):

You know the answer to #7?

OpenStudy (anonymous):

no..

OpenStudy (ranga):

The equations that you found for each segment: Launch point to A, A to B and B to C are applicable only to those segments and not beyond. The coordinates were: Launch point (1,2) A is (2,3); B is (3,5); C is (4,9) So the equation for the launch point to point A applies only to that line segment. The equation is no longer applicable for the trajectory of the vehicle past those points. The domain and range of the first equation is: domain [1,2] and range [2,3]. Similar restrictions for the other two equations.

OpenStudy (anonymous):

thats the answer?^

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

#8?

OpenStudy (ranga):

Given any function, the domain is all the allowed values for x. The range is all the values y can take. The function is not defined for x values not in the domain. Therefore we cannot use the equation for x values not in the domain because the equation is no longer valid. The range gives the minimum y value to the maximum y value. For this problem where we are dealing with 3 straight lines, I don't think we have to worry about division by 0, etc. which is one of the reasons why some x values may be excluded from the domain. Here, the restriction in the domain comes because the equation of the line is applicable only to each segment and not beyond.

OpenStudy (anonymous):

answer^

OpenStudy (anonymous):

?

OpenStudy (ranga):

Yes. But my answers to #7 & #8 are mixed in with some information to you. Hopefully you can separate them.

OpenStudy (anonymous):

thank you so much!

OpenStudy (ranga):

You are welcome.

OpenStudy (ranga):

Once again, sorry to hear about your loss.